OpenStudy (anonymous):

What is the equation of the line that has a slope of -1/3 and a y-intercept 5/2?

OpenStudy (anonymous):

Plug it into the slope-intercept form. y = mx + b Where m is the slope, and b is the y-intercept.

OpenStudy (anonymous):

What do you get?

OpenStudy (anonymous):

i can answer this.hold on

OpenStudy (anonymous):

@ACTIONSPORTS90210 Please don't give answers, we are here to guide the asker to the solution, so that they get it on their own.

OpenStudy (anonymous):

· Find the slope and the y-intercept, and write an equation of the line. · Given the slope and a point on the line, write an equation of the line. · Given two points, write the equation of a line.

OpenStudy (anonymous):

follow that first.and i wont @iGreen.

OpenStudy (anonymous):

...... What?

OpenStudy (anonymous):

@ms1231988 Can you make the equation now?

OpenStudy (anonymous):

Y=MX+B So 5/2=1/3+b?

OpenStudy (anonymous):

Im so confused, im sorry.

OpenStudy (anonymous):

No, you keep 'y' the same..and you keep 'x' there.. \(\sf y = mx + b\) We are going to replace \(\sf m\) with \(\sf -1/3\) and \(\sf b\) with \(\sf 5/2\).

OpenStudy (anonymous):

y=-1/3+5/2 ...?

OpenStudy (anonymous):

Yes..but we keep the 'x' there: \(\sf y = -1/3x + 5/2\)

OpenStudy (anonymous):

Thanks